### v1

@@ 2.4 Proportional Control Methodology
Time step, $\mathrm{dt}$ 0.1
@@ 3.1 Baselines
Figure 4: Temperature fields of the baselines shown in the mid-plane cross-section ($y=0$) of the domain for (a) $\mathrm{Ra}=500$ at $t=400$ and (b) for $\mathrm{Ra}=750$ for class 3 (run 5) at time $t=5000$. Temperature fields of runs 1–4 of the other classes of baselines for $\mathrm{Ra}=750$ are shown in Appendix C. Video files are provided in Online Resource 1 and 2 for $\mathrm{Ra}=500$ and 750 respectively (see Appendix A).
In Fig. 4, we plot the shapes of the convection rolls at the final time step of the baselines. Although training for $\mathrm{Ra}=750$ is performed with all 5 baseline runs, for brevity, Fig. 4(b) shows the topology for class 3 (run 5) only, since this is the baseline from which the deterministic agent evaluation and proportional control cases are simulated. The topologies of the remaining baseline runs for $\mathrm{Ra}=750$ are shown in Fig. 20 of Appendix C. At the warmest regions, the $v$ velocities are positive and maximum indicating upward flow, and in the coolest regions, $v$ is negative and minimum indicating flow towards the bottom wall. We define the length scale of the convection rolls as the distance between the warm and cold regions adjacent to each other at points where the temperatures are the maximum and minimum, respectively (i.e., $v$ is maximum and minimum). In both cases, the convection rolls have a length scale of approximately $\pi$, an observation in line with results of linear stability analysis [4]. Note that videos of the temperature field evolution for these baselines are provided in Online Resource 1 for $\mathrm{Ra}=500$ and Online Resource 2 for $\mathrm{Ra}=750$. Visible from the fields, all baseline shapes for $\mathrm{Ra}=500$ and 750 are irregularly shaped, with no particular spatial pattern. Class 2 and class 3 baselines for $\mathrm{Ra}=750$ exhibit spatial oscillations about a mean shape.
@@ 3.3.2 $\mathrm{Ra}=750$
To better understand the learned policy, we observe the plots in 11(c)–(f), where four segments are chosen at random to analyse the action chosen based on the observed states. Opposition control-like relationships are visible in segments 1 and 20, where from $t=6000-7000$, high temperature observations at the mid-plane cause the corresponding actuations to be lower than $T_{H,0}=2$. However, in segments 40 and 56, an opposite trend is visible. In segment 40 at $t=6500$, a decrease in $T_{\mathrm{obs}}$ is accompanied by a decrease in the actuation temperatures as well, and in segment 56 at $t=6500$, an increase in $T_{\mathrm{obs}}$ is accompanied by an increase in $T_{\mathrm{act}}$. Similar to $\mathrm{Ra}=500$, the delays between the increase/decrease of $T_{\mathrm{obs}}$ and increase/decrease of $T_{\mathrm{act}}$ are segment-dependent and not consistent across segments, which differentiates the agent-learned control policy from a simple opposition control policy. The qualitative results of the learned control strategy are similar to that observed in $\mathrm{Ra}=500$. The fact that such differences in the policies are observed, while the MARL setup implies that the DRL agent is shared across segments, indicates that complex non-linear strategies that depend on the state in a complex fashion have been chosen by the agent.
@@ 4 Conclusion
These results open opportunities for a number of future investigations. Various configurations of agents can be adopted for the control of RBC at larger domains and larger $\mathrm{Ra}$. For example, instead of a single agent having as input the states corresponding to a single segment, a set of multiple neighbouring state segments can be assigned as the input state to a single agent. The input states would then contain information from a larger window of the domain that is influenced by the action imposed in the control segment. We are also working on investigating how the results presented here can be extended to higher $\mathrm{Ra}$ cases, in which convection is both stronger and more chaotic.
@@
Acknowledgements Part of the deep-learning-model training was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS). Ricardo Vinuesa acknowledges financial support from ERC grant no. ‘2021-CoG-101043998, DEEPCONTROL’. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@
Appendix A Supplementary information
@@
Appendix B Open source code release on Github
All the codes, scripts, and post-processing tools used in this work are made available on Github together with readmes and user instructions. Reasonable user support will be provided through the issue tracker of the corresponding Github repository. The url for the repository is https://github.com/KTH-FlowAI/DRL_MARL_RayleighBenard3D_Control.
@@ Appendix B Open source code release on Github
Appendix C Convection structure for $\mathrm{Ra}=750$ baselines

### v2

@@ 2.4 Proportional Control Methodology
Time step, $\mathrm{\Delta t}$ 0.1
@@ 3.1 Baselines
Figure 4: Temperature fields of the baselines shown in the mid-plane cross-section ($y=0$) of the domain for (a) $\mathrm{Ra}=500$ at $t=400$ and (b) for $\mathrm{Ra}=750$ for class 3 (run 5) at time $t=5000$. Temperature fields of runs 1–4 of the other classes of baselines for $\mathrm{Ra}=750$ are shown in Appendix D. Video files are provided in Online Resource 1 and 2 for $\mathrm{Ra}=500$ and 750 respectively (see Appendix B).
In Fig. 4, we plot the shapes of the convection rolls at the final time step of the baselines. Although training for $\mathrm{Ra}=750$ is performed with all 5 baseline runs, for brevity, Fig. 4(b) shows the topology for class 3 (run 5) only, since this is the baseline from which the deterministic agent evaluation and proportional control cases are simulated. The topologies of the remaining baseline runs for $\mathrm{Ra}=750$ are shown in Fig. 22 of Appendix D. At the warmest regions, the $v$ velocities are positive and maximum indicating upward flow, and in the coolest regions, $v$ is negative and minimum indicating flow towards the bottom wall. We define the length scale of the convection rolls as the distance between the warm and cold regions adjacent to each other at points where the temperatures are the maximum and minimum, respectively (i.e., $v$ is maximum and minimum). In both cases, the convection rolls have a length scale of approximately $\pi$, an observation in line with results of linear stability analysis [4]. We perform a convergence analysis with different simulation time steps $\mathrm{\Delta t}$ to reinforce this result, which further justifies our choice of $\mathrm{\Delta t}=0.1$. The analysis shows that simulations with $\mathrm{\Delta t}\leq 0.5$ are reasonable choices of time steps to obtain solutions with length scales of convection rolls that agree with linear stability analysis. Details of the convergence analysis are provided in Appendix A.
Note that videos of the temperature field evolution for these baselines are provided in Online Resource 1 for $\mathrm{Ra}=500$ and Online Resource 2 for $\mathrm{Ra}=750$. Visible from the fields, all baseline shapes for $\mathrm{Ra}=500$ and 750 are irregularly shaped, with no particular spatial pattern. Class 2 and class 3 baselines for $\mathrm{Ra}=750$ exhibit spatial oscillations about a mean shape.
@@ 3.3.2 $\mathrm{Ra}=750$
To better understand the learned policy, we observe the plots in 11(c)–(e), where four segments are chosen at random to analyse the action chosen based on the observed states. Opposition control-like relationships are visible in segments 1 and 20, where from $t=6000-7000$, high temperature observations at the mid-plane cause the corresponding actuations to be lower than $T_{H,0}=2$. However, in segments 40 and 56, an opposite trend is visible. In segment 40 at $t=6500$, a decrease in $T_{\mathrm{obs}}$ is accompanied by a decrease in the actuation temperatures as well, and in segment 56 at $t=6500$, an increase in $T_{\mathrm{obs}}$ is accompanied by an increase in $T_{\mathrm{act}}$. Similar to $\mathrm{Ra}=500$, the delays between the increase/decrease of $T_{\mathrm{obs}}$ and increase/decrease of $T_{\mathrm{act}}$ are segment-dependent and not consistent across segments, which differentiates the agent-learned control policy from a simple opposition control policy. The qualitative results of the learned control strategy are similar to that observed in $\mathrm{Ra}=500$. The fact that such differences in the policies are observed, while the MARL setup implies that the DRL agent is shared across segments, indicates that complex non-linear strategies that depend on the state in a complex fashion have been chosen by the agent.
@@ 4 Conclusion
These results open opportunities for a number of future investigations. Various configurations of agents can be adopted for the control of RBC at larger domains and larger $\mathrm{Ra}$. For example, instead of a single agent having as input the states corresponding to a single segment, a set of multiple neighbouring state segments can be assigned as the input state to a single agent. The input states would then contain information from a larger window of the domain that is influenced by the action imposed in the control segment. We are also working on investigating how the results presented here can be extended to higher $\mathrm{Ra}$ cases, in which convection is both stronger and more chaotic. In other flow-control problems, DRL-based control has shown promising results in the turbulent regimes, such as 3D cylinders [55], so one could also expect positive results for higher $\mathrm{Ra}$ number cases.
DRL-based control has shown promising results in the turbulent regimes for a 3D cylinder, as we see in e.g. flow control for the wake of 3D cylinder (Suarez et al. 2024)
We believe that the present work is an important milestone that takes DRL from simple 2D cases [63] to complex non-stationary 3D cases. In the future, we will push this further to consider fully turbulent 3D cases, which are relevant for many industrial applications, such as chemical reactors, material manufacturing, and thermal energy systems, where maintaining stable temperature gradients and flow patterns is crucial in order to prevent undesirable instabilities and to enhance efficiency. In other examples such as thermal insulation of buildings, understanding the principles of control in RBC can be crucial to minimize the energy losses of the heating or cooling system.
@@
Acknowledgements The authors extend their acknowledgement to the National Academic Infrastructure for Supercomputing in Sweden (NAISS) for resources provided for this study, and to the European Research Council (ERC) for financial support. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
@@
[…]
Funding Information Part of the deep-learning-model training was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS). Ricardo Vinuesa acknowledges financial support from ERC grant no. ‘2021-CoG-101043998, DEEPCONTROL’.
@@
Appendix A Convergence Analysis
In this section, we simulate the baseline for $Ra=500$ with different time steps Δt=0.01, 0.1, $\mathrm{\Delta t}=0.01,\ 0.1,\ 0.2$ and $0.5$ in order to justify our choice of $\mathrm{\Delta t}=0.1$. For each $\mathrm{\Delta t}$, we run five simulations with the same parameters. We note here that for $\mathrm{\Delta t}>0.5$, the solution is found to be numerically unstable and the values of Nu obtained are Nan. Shown in Fig. 20 are the resulting Nu evolutions for $\mathrm{\Delta t}\leq 0.5$. We see that in each case, the Nu can reach either a steady or unsteady solution with small fluctuations about the mean. We also note that the mean Nu (averaged from $t=400-1000$) for each of the five simulations for each $\mathrm{\Delta t}$, lies within the range of 1.23-1.27.
We also observe the flow structures in the temperature fields shown in Fig. 21. The temperature fields for each $\mathrm{\Delta t}$ are from a single randomly selected simulation out of the five solutions, at the end time $t=1000$. By visual inspection, we note that the convection rolls have a length scale of $\pi$ that is roughly uniform along the axis of the rolls. Thus, solutions with $\mathrm{\Delta t}\leq 0.5$ qualitatively have the same spatial and temporal features, and we deem any of them fit to use as baselines for control.
Figure 20: Time evolutions of Nu with various time steps $\mathrm{\Delta t}$. For each $\mathrm{\Delta t}$, five simulations with the same parameters are shown.
Figure 21: Mid-plane temperature fields at $t=1000$ for the simulations with different time steps as shown in Fig. 20.
The thermal relaxation time for the non-dimensional Rayleigh-Bénard problem is $\sqrt{\mathrm{Ra}\cdot\mathrm{Pr}}$ [4] which for $Ra=500$ and $\mathrm{Pr}=0.7$ is $\approx 18.7$ non-dimensional time units. A motivating factor for selecting $\mathrm{\Delta t}=0.1$ (instead of 0.01 or any value $\leq 0.5$) for the remainder of the paper is that this is small enough to be at least 1/100th the thermal relaxation time, and large enough to ensure quicker simulations. Furthermore, during the control simulations, a low $\mathrm{\Delta t}$ ensures that any numerical instabilities that may arise due to the sharp changes in temperature actuations at the bottom wall can be prevented.
Appendix B Supplementary information
@@
Appendix C Open source code release on Github
All the codes, scripts, and post-processing tools used in this work will be made available on Github together with readmes and user instructions, over publication of this manuscript. Reasonable user support will be provided through the issue tracker of the corresponding Github repository. The url for the repository is https://github.com/KTH-FlowAI/DRL_MARL_RayleighBenard3D_Control.
@@ Appendix C Open source code release on Github
Appendix D Convection structure for $\mathrm{Ra}=750$ baselines